OpenStudy (anonymous):

(b/c)=7, how do you solve for c?

OpenStudy (anonymous):

Make up a number for b and divide it by 7

OpenStudy (anonymous):

i'm using whatever c equals to solve an equation, so that doesn't work, thanks though!

OpenStudy (campbell_st):

this is change of subject... multiply both sides of the equation by c which gives b = 7c now divide both sides by 7 so c = b/7 question done.

OpenStudy (anonymous):

so how would you solve: (a+b)=3 (b+c)=7 and you need to solve (a+b)/(b+c)

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

so its simply 3/7 since (a+b) = 3 ---------- (b + c) = 7

OpenStudy (anonymous):

@campbell_st I don' think that is right, the question is

OpenStudy (anonymous):

if a, b and c are real numbers such that (a/b)=3 and (b/c)=7, then (a+b)/(b+c) equals what?

OpenStudy (anonymous):

@campbell_st i wrote the wrong problem before ii am sorry

OpenStudy (campbell_st):

well you can look at it as from the 1st equation then b = a/3 from the 2nd b = 7c so a/3 = 7c or a = 21c so a + b = 21c + 7c b + c = 7c + c so (a +b)/(b+c) = 28c/8c = 3.5